OpenStudy (anonymous):

what is (5/9) + (1/12) ?

OpenStudy (anonymous):

you would have to find common denominators for 9 and 12 , which would be 36, so you would multiply 5/9 by 4 and multiply 1/12 by 3 , so it would be 20/36+3/36 , only add the numerator which is 23, so the answer is 23/36 , since it cannot be simplified , you leave it like that ..

OpenStudy (anonymous):

To get this you need the common denominator. This means denominators of both must be equal in value to each other. This is easily done by multiplying each fraction by the denominator of the other: (5*12)/(9*12) for the fraction 5/9 (1*9)/(12*9) for the fraction 1/12 this gives 5/9 = 60/108 and 1/12 = 9/108 As you only need to add the numerators the question becomes 60 + 9 = 69 Accounting for the numerators again the answer is 69/108. However this can be simplified by dividing by three to give (69/3)/(108/3)= 23/36. As you can't simplify farther the answer is 5/9 + 1/12 = 23/36